OpenStudy (anonymous):

Evaluate the expression. r = <5, 9, -4>, v = <3, 3, -8>, w = <9, -6, -1> v ⋅ w explain?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

they are asking for the dot product right?

OpenStudy (anonymous):

yes , but how do you do the dot product

OpenStudy (anonymous):

and Hi!

OpenStudy (misty1212):

\[<x_1,x_2,x_3>\cdot <y_1,y_2,y_3>=x_1\times y_1+x_2\times y_2+x_3\times y_3\] i.e. multiply and add that is all

OpenStudy (anonymous):

so its 17!

OpenStudy (misty1212):

\[27-18+8\]

OpenStudy (misty1212):

yeah